OpenStudy (arbyj):

helper will get a medal!!! the height of cylinder is 12 and the diameter is 14. what is the surface area? I got 835.7 but, I'm not sure if it's correct

OpenStudy (anonymous):

835.664

OpenStudy (arbyj):

@CheshireCat how bout the volume?

OpenStudy (anonymous):

1847.26

OpenStudy (anonymous):

so

OpenStudy (arbyj):

what would the equation look like? @CheshireCat

OpenStudy (whpalmer4):

a cylinder has two ends, which are circles, and the side, which is a rectangle with the opposite edges joined. Area of a circle given the diameter is \[A = \pi r^2 = \pi (\frac{d}{2})^2 = \frac{\pi d^2} {4}\] You'll have two of those, one for each end. The rectangle that is rolled to make the side of the cylinder has the same height as the cylinder, and its width is the circumference of the end, or \(C = \pi d\). Therefore, the area of the side of the cylinder is \[A = l*w = h*C = \pi h d\] The surface area of the cylinder will be the area of each of the ends plus the side. You should be able to figure out the formula from the pieces I gave you. The number you computed for surface area is correct. For volume of the cylinder, it is simply the area of the end multiplied by the height: \[V = \pi r^2 h = \frac{\pi d^2 h}{4}\]